OpenStudy (anonymous):

At which value will the graph of y=csc x have a zero? A, pi/2 B. pi/4 C. pi/3 D. none of the above are correct.

jimthompson5910 (jim_thompson5910):

y = csc(x) y = 1/sin(x) 0 = 1/sin(x) 0*sin(x) = 1 0 = 1 Since this equation is never true, there are no zeros for the function

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Agreed thanks a lot i was just a bit unsure